Exercise 7-4 Your accounts receivable clerk, Mitra Adams, to whom you pay a salary of $2,265 per month, has just purchased a new Acura. You decide to test the accuracy of the accounts receivable balance of $123,820 as shown in the ledger. The following information is available for your first year in business. (1) Collections from customers $298,980 (2) Merchandise purchased 483,200 (3) Ending merchandise inventory 135,900 (4) Goods are marked to sell at 40% above cost Compute an estimate of the ending balance of accounts receivable from customers that should appear in the ledger and any apparent shortages. Assume that all sales are made on account. The ending balance of accounts receivable from customers $ Apparent shortage $
Answer #1

Answer:

Since it is first year of operation:

Beginning merchandise inventory = 0

Cost of goods sold = Merchandise purchased - Ending merchandise inventory = 483200 - 135900 =$347,300

On account Sales = 347300 * (1 + 40%) = $486,220

Collections = $298,980

Estimated ending balance of accounts receivable = 486220 - 298980 = $187,240

Apparent shortage = 187240 -123820 = $63,420

Hence:

Estimated ending balance of accounts receivable = $187,240

Apparent shortage = $63,420